# Model-independent predictions for decays of double-heavy hadrons into pairs of heavy hadrons

### v2

@@ II.3 Pair of Heavy Hadrons
Let us now consider a pair of heavy hadrons at the positions $+\frac{1}{2}\bm{r}$ and $-\frac{1}{2}\bm{r}$. We take the first heavy hadron to contain the heavy quark $Q$ and the second to contain either the heavy quark $Q$ or the heavy antiquark $\bar{Q}$.
The heavy-hadron state in Eq. (3) is expressed in terms of the direct product of a spin/parity state $\frac{1}{2}^{+}$ for a heavy quark $Q$ (or $\frac{1}{2}^{-}$ for a heavy antiquark $\bar{Q}$) and an angular-momentum/parity state $j^{\pi}$ for a static hadron. A heavy-hadron-pair state can be expressed in terms of the direct product of two spin/parity states for $Q$ and $Q$ or for $Q$ and $\bar{Q}$ and a static-hadron-pair state of the form in Eq. (5). The angular-momentum operator that generates rotations of the light-QCD fields and the heavy-quark and heavy-antiquark spins is the static angular-momentum vector
@@
where $\bm{S}_{Q}$ is the sum of the two spin vectors for $Q$ and $Q$ or for $Q$ and $\bar{Q}$. We denote the quantum numbers for the total-heavy-spin vector $\bm{S}_{Q}$ by $(S_{Q},m)$. For a heavy-hadron pair, the static-angular-momentum vector coincides with the sum of the spin vectors for the two heavy hadrons, $\bm{J}_{\textup{static}}=\bm{J}_{1}+\bm{J}_{2}$. Therefore, we denote the quantum numbers for the static angular momentum vector $\bm{J}_{\textup{static}}$ by $(S,m_{S})$.
The spin/parity state corresponding to $Q$ at the position $+\frac{1}{2}\bm{r}$ is $\frac{1}{2}^{+}(+)$. The spin/parity state corresponding to $Q$ (or $\bar{Q}$) at the position $-\frac{1}{2}\bm{r}$ is $\frac{1}{2}^{+}(-)$ (or $\frac{1}{2}^{-}(-)$). Their direct product can be decomposed into states with total-heavy-spin quantum numbers $(S_{Q},m)$:
$\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{\pm}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}=\sum_{m^{\prime},m^{\prime\prime}}\biggl{\langle}\begin{matrix}\tfrac{1}{2}&\tfrac{1}{2}\\[-5.0pt] m^{\prime}&m^{\prime\prime}\end{matrix}\bigg{|}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}\biggl{|}\begin{matrix}\tfrac{1}{2}^{+}\\[-5.0pt] m^{\prime}\end{matrix}(+)\biggr{\rangle}\biggl{|}\begin{matrix}\tfrac{1}{2}^{\pm}\\[-5.0pt] m^{\prime\prime}\end{matrix}(-)\biggr{\rangle},$ (15)
with the two choices for the parity superscript $\pm$ corresponding to $Q$ or $\bar{Q}$ at the position $-\frac{1}{2}\bm{r}$. We have adopted the standard order of writing the (+)(+)( + ) state first and the ()(-)( - ) state second. Reversal of this order produces a sign:
$\biggl{|}\Bigl{(}\tfrac{1}{2}^{\pm}(-),\tfrac{1}{2}^{+}(+)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}=(-1)^{S_{Q}}\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{\pm}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}.$ (16)
The sign is the product of a factor $(-1)^{S_{Q}-1}$ from the symmetries of Clebsch-Gordan coefficients and a factor $-1$ from changing the order of the fermionic operators for the heavy quark and (anti)quark. From Eqs. (II.1), the transformations under $\mathcal{P}$ and $\mathcal{C}$ of the total-heavy-spin state in Eq. (15) are:
$\mathcal{P}\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{\pm}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}=$$\pm\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(-),\tfrac{1}{2}^{\pm}(+)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle},$ (17e)
$\mathcal{C}\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{\pm}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}=$ $\biggl{|}\Bigl{(}\tfrac{1}{2}^{-}(+),\tfrac{1}{2}^{\mp}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}.$ (17j)
On the right side of Eq. (17e), the position labels ()(-)( - ) and (+)(+)( + ) can be put into the standard order by using Eq. (16). Thus, Eqs. (16) and (II.3) imply that the state in Eq. (15) is an eigenstate of $[\mathcal{C}]\mathcal{P}$ with eigenvalue ±$\pm(-1)^{S_{Q}}$.
@@
$\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{\pm}(-)\Bigr{)}S_{Q},\Bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\Bigr{)}j^{\prime}\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}\\ =\sum_{m,\lambda}\biggl{\langle}\begin{matrix}S_{Q}&j^{\prime}\\[-5.0pt] m&\lambda\end{matrix}\bigg{|}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{\pm}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}\biggl{|}\Bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\Bigr{)}\begin{matrix}j^{\prime}\\[-5.0pt] \lambda\end{matrix}\biggr{\rangle}.$ (18)
An alternative basis for heavy-hadron-pair states that is often more convenient can be obtained by starting from direct products of a heavy-hadron state $\bigl{(}\frac{1}{2}^{+},j_{1}^{\pi_{1}}\bigr{)}J_{1}(+)$ at the position $+\frac{1}{2}\bm{r}$ and a heavy-hadron state $\bigl{(}\frac{1}{2}^{\pm},j_{2}^{\pi_{2}}\bigr{)}J_{2}(-)$ at the position $-\frac{1}{2}\bm{r}$, with the heavy-hadron states defined in Eq. (3). The direct product can be decomposed into states with static-angular-momentum quantum numbers $(S,m_{S})$:
$\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}\\ =\sum_{M_{1},M_{2}}\biggl{\langle}\begin{matrix}J_{1}&J_{2}\\[-5.0pt] M_{1}&M_{2}\end{matrix}\bigg{|}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}\biggl{|}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}\begin{matrix}J_{1}\\[-5.0pt] M_{1}\end{matrix}(+)\biggr{\rangle}\biggl{|}\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}\Bigr{)}\begin{matrix}J_{2}\\[-5.0pt] M_{2}\end{matrix}(-)\biggr{\rangle},$ (19)
@@
$\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-),\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}\\ =(-1)^{4J_{1}J_{2}+J_{1}+J_{2}-S}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}.$ (20)
@@
The sign is the product of a factor $(-1)^{J_{1}+J_{2}-S}$ from the symmetries of Clebsch-Gordan coefficients and a factor $(-1)^{4J_{1}J_{2}}$ from changing the order of the operators for the heavy hadrons $(\frac{1}{2}^{+},j_{1}^{\pi_{1}})J_{1}$ and $(\frac{1}{2}^{\pm},j_{2}^{\pi_{2}})J_{2}$. [^3] The factor $(-1)^{4J_{1}J_{2}}$ is $-1$ if both heavy hadrons are baryons and $+1$ otherwise.
@@
From Eqs. (II.1), the transformation properties under $\mathcal{P}$ and $\mathcal{C}$ of the hadron-pair state in Eq. (19) are:
@@
$\mathcal{P}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}$ $\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}$ (21c)
@@
=±$=\pm\pi_{1}\pi_{2}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(-),\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(+)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle},$ (21f)
@@
$\mathcal{C}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}$ $\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}$ (21i)
@@
$=\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{-},j_{1}^{\pi_{1}(-1)^{2j_{1}}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{\mp},j_{2}^{\pi_{2}(-1)^{2j_{2}}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}.$ (21l)
@@
On the right side of Eq. (21f), the position labels ()(-)( - ) and (+)(+)( + ) can be put into the standard order by using Eq. (20).
@@
There are special cases in which the hadron-pair state in Eq. (19) is a simultaneous eigenstate of $[\mathcal{C}]\mathcal{P}$ with eigenvalue $\eta^{\prime}$. For $(\bm{3},\bm{3})$ sources, the hadron-pair state is a simultaneous eigenstate of $\mathcal{P}$ when the two heavy hadrons are identical, which requires $J_{2}=J_{1}$ and $j_{2}^{\pi_{2}}=j_{1}^{\pi_{1}}$. After reordering the ()(-)( - ) and (+)(+)( + ) states using Eq. (20), the sign is the product of $\pi_{1}^{2}=1$ and $(-1)^{4J_{1}^{2}+2J_{1}-S}$. For $(\bm{3},\bm{3}^{\ast})$ sources, the hadron-pair state is a simultaneous eigenstate of $\mathcal{C}\mathcal{P}$ when the two heavy hadrons are charge conjugates, which requires $J_{2}=J_{1}$ and $j_{2}^{\pi_{2}}=j_{1}^{\pi_{1}(-1)^{2j_{1}}}$. After reordering the ()(-)( - ) and (+)(+)( + ) states using Eq. (20), the sign is the product of $-\pi_{1}^{2}(-1)^{2j_{1}}=(-1)^{2J_{1}}$ and $(-1)^{4J_{1}^{2}+2J_{1}-S}$. Since $2J_{1}(2J_{1}+1)$ is even, the $[\mathcal{C}]\mathcal{P}$ eigenvalue reduces to
$\eta^{\prime}=\begin{cases}(-1)^{S}&\text{for $(\bm{3},\bm{3})$ sources,}\\ (-1)^{S+2J_{1}}&\text{for $(\bm{3},\bm{3}^{\ast})$ sources.}\end{cases}$ (22)
@@
Note that $S$ is an integer if the hadron-pair state is an eigenstate of $[\mathcal{C}]\mathcal{P}$. Outside of these special cases, the hadron-pair state in Eq. (19) is an equal-amplitude superposition of $\eta^{\prime}=+1$ and $\eta^{\prime}=-1$. In this case, the normalized projection onto each value of $\eta^{\prime}$ is
@@
$\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix},\eta^{\prime}\biggr{\rangle}\\ \equiv\sqrt{2}\,\Pi_{\eta^{\prime}}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle},$ (23)
@@
$\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}=(-1)^{2j_{1}}\sqrt{\tilde{J}_{1}\tilde{J}_{2}}\\ \times\sum_{j^{\prime},S_{Q}}\sqrt{\tilde{\jmath}^{\prime}\tilde{S}_{Q}}\begin{Bmatrix}\tfrac{1}{2}&\tfrac{1}{2}&S_{Q}\\ j_{1}&j_{2}&j^{\prime}\\ J_{1}&J_{2}&S\\ \end{Bmatrix}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{\pm}(-)\Bigr{)}S_{Q},\Bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\Bigr{)}j^{\prime}\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle},$ (24)
where $\tilde{J}=2J+1$ and $\left\{\begin{smallmatrix}j_{1}&j_{2}&j_{3}\\ j_{4}&j_{5}&j_{6}\\ j_{7}&j_{8}&j_{9}\end{smallmatrix}\right\}$ is a Wigner 9-$j$ symbol. The sign $(-1)^{2j_{1}}$ is due to changing the order of the light-QCD operator for the static hadron $j_{1}^{\pi_{1}}$ and the fermionic operator for the heavy quark or antiquark state $\tfrac{1}{2}^{\pm}$.
III Decays of a Double-Heavy Hadron
In this section, we determine the general form of the coupling potential between a confining B-O potential and a meson-pair potential. We use it to derive selection rules for decays of a double-heavy hadron and analytic results for relative partial decay rates.

### v3

@@ II.3 Pair of Heavy Hadrons
Let us now consider a pair of heavy hadrons at the positions $+\frac{1}{2}\bm{r}$ and $-\frac{1}{2}\bm{r}$. We take the first heavy hadron to contain the heavy quark $Q$ and the second to contain the heavy antiquark $\bar{Q}$.
The heavy-hadron state in Eq. (3) is expressed in terms of the direct product of a spin/parity state $\frac{1}{2}^{+}$ for a heavy quark $Q$ (or $\frac{1}{2}^{-}$ for a heavy antiquark $\bar{Q}$) and an angular-momentum/parity state $j^{\pi}$ for a static hadron. A heavy-hadron-pair state can be expressed in terms of the direct product of two spin/parity states for $Q$ and $\bar{Q}$ and a static-hadron-pair state of the form in Eq. (5). The angular-momentum operator that generates rotations of the light-QCD fields and the heavy-quark and heavy-antiquark spins is the static angular-momentum vector
@@
where $\bm{S}_{Q}$ is the sum of the two spin vectors for $Q$ and $\bar{Q}$. We denote the quantum numbers for the total-heavy-spin vector $\bm{S}_{Q}$ by $(S_{Q},m)$. For a heavy-hadron pair, the static-angular-momentum vector coincides with the sum of the spin vectors for the two heavy hadrons, $\bm{J}_{\textup{static}}=\bm{J}_{1}+\bm{J}_{2}$. Therefore, we denote the quantum numbers for the static angular momentum vector $\bm{J}_{\textup{static}}$ by $(S,m_{S})$.
The spin/parity state corresponding to $Q$ at the position $+\frac{1}{2}\bm{r}$ is $\frac{1}{2}^{+}(+)$. The spin/parity state corresponding to $\bar{Q}$ at the position $-\frac{1}{2}\bm{r}$ is $\frac{1}{2}^{-}(-)$. Their direct product can be decomposed into states with total-heavy-spin quantum numbers $(S_{Q},m)$:
$\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{-}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}=\sum_{m^{\prime},m^{\prime\prime}}\biggl{\langle}\begin{matrix}\tfrac{1}{2}&\tfrac{1}{2}\\[-5.0pt] m^{\prime}&m^{\prime\prime}\end{matrix}\bigg{|}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}\biggl{|}\begin{matrix}\tfrac{1}{2}^{+}\\[-5.0pt] m^{\prime}\end{matrix}(+)\biggr{\rangle}\biggl{|}\begin{matrix}\tfrac{1}{2}^{-}\\[-5.0pt] m^{\prime\prime}\end{matrix}(-)\biggr{\rangle}.$ (15)
$\biggl{|}\Bigl{(}\tfrac{1}{2}^{-}(-),\tfrac{1}{2}^{+}(+)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}=(-1)^{S_{Q}}\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{-}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}.$ (16)
The sign is the product of a factor $(-1)^{S_{Q}-1}$ from the symmetries of Clebsch-Gordan coefficients and a factor $-1$ from changing the order of the fermionic operators for the heavy quark and antiquark. From Eqs. (1), the transformations under $\mathcal{P}$ and $\mathcal{C}$ of the total-heavy-spin state in Eq. (15) are:
$\mathcal{P}\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{-}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}=$ $-\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(-),\tfrac{1}{2}^{-}(+)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle},$ (17a)
$\mathcal{C}\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{-}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}=$ $\biggl{|}\Bigl{(}\tfrac{1}{2}^{-}(+),\tfrac{1}{2}^{+}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}.$ (17b)
On the right side of Eq. (17a), the position labels ()(-)( - ) and (+)(+)( + ) can be put into the standard order by using Eq. (16). Thus, Eqs. (16) and (17) imply that the state in Eq. (15) is an eigenstate of $\mathcal{C}\mathcal{P}$ with eigenvalue $(-1)^{S_{Q}+1}$.
@@
$\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{-}(-)\Bigr{)}S_{Q},\Bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\Bigr{)}j^{\prime}\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}\\ =\sum_{m,\lambda}\biggl{\langle}\begin{matrix}S_{Q}&j^{\prime}\\[-5.0pt] m&\lambda\end{matrix}\bigg{|}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}\biggl{|}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{-}(-)\Bigr{)}\begin{matrix}S_{Q}\\[-5.0pt] m\end{matrix}\biggr{\rangle}\biggl{|}\Bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\Bigr{)}\begin{matrix}j^{\prime}\\[-5.0pt] \lambda\end{matrix}\biggr{\rangle}.$ (18)
An alternative basis for heavy-hadron-pair states that is often more convenient can be obtained by starting from direct products of a heavy-hadron state $\bigl{(}\frac{1}{2}^{+},j_{1}^{\pi_{1}}\bigr{)}J_{1}(+)$ at the position $+\frac{1}{2}\bm{r}$ and a heavy-hadron state $\bigl{(}\frac{1}{2}^{-},j_{2}^{\pi_{2}}\bigr{)}J_{2}(-)$ at the position $-\frac{1}{2}\bm{r}$, with the heavy-hadron states defined in Eq. (3). The direct product can be decomposed into states with static-angular-momentum quantum numbers $(S,m_{S})$:
$\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}\\ =\sum_{M_{1},M_{2}}\biggl{\langle}\begin{matrix}J_{1}&J_{2}\\[-5.0pt] M_{1}&M_{2}\end{matrix}\bigg{|}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}\biggl{|}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}\begin{matrix}J_{1}\\[-5.0pt] M_{1}\end{matrix}(+)\biggr{\rangle}\biggl{|}\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}\Bigr{)}\begin{matrix}J_{2}\\[-5.0pt] M_{2}\end{matrix}(-)\biggr{\rangle},$ (19)
@@
$\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-),\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}\\ =(-1)^{4J_{1}J_{2}+J_{1}+J_{2}-S}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}.$ (20)
@@
The sign is the product of a factor $(-1)^{J_{1}+J_{2}-S}$ from the symmetries of Clebsch-Gordan coefficients and a factor $(-1)^{4J_{1}J_{2}}$ from changing the order of the operators for the heavy hadrons $(\frac{1}{2}^{+},j_{1}^{\pi_{1}})J_{1}$ and $(\frac{1}{2}^{-},j_{2}^{\pi_{2}})J_{2}$. [^2] The factor $(-1)^{4J_{1}J_{2}}$ is $-1$ if both heavy hadrons are baryons and $+1$ otherwise.
@@
From Eqs. (4), the transformation properties under $\mathcal{P}$ and $\mathcal{C}$ of the hadron-pair state in Eq. (19) are:
@@
$\mathcal{P}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}$ $\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}$
@@
$=-\pi_{1}\pi_{2}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(-),\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(+)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle},$ (21a)
@@
$\mathcal{C}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}$ $\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}$
@@
$=\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{-},j_{1}^{\pi_{1}(-1)^{2j_{1}}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{+},j_{2}^{\pi_{2}(-1)^{2j_{2}}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}.$ (21b)
@@
On the right side of Eq. (21a), the position labels ()(-)( - ) and (+)(+)( + ) can be put into the standard order by using Eq. (20).
@@
The hadron-pair state in Eq. (19) is a simultaneous eigenstate of $\mathcal{C}\mathcal{P}$ when the two heavy hadrons are charge conjugates, which requires $J_{2}=J_{1}$ and $j_{2}^{\pi_{2}}=j_{1}^{\pi_{1}(-1)^{2j_{1}}}$. After reordering the ()(-)( - ) and (+)(+)( + ) states using Eq. (20), the sign is the product of $-\pi_{1}^{2}(-1)^{2j_{1}}=(-1)^{2J_{1}}$ and $(-1)^{4J_{1}^{2}+2J_{1}-S}$. Since $2J_{1}(2J_{1}+1)$ is even, the $\mathcal{C}\mathcal{P}$ eigenvalue reduces to
$\eta^{\prime}=(-1)^{S+2J_{1}},$ (22)
@@
where we have used the fact that $S$ is an integer since $J_{1}=J_{2}$. Outside of this special case, the hadron-pair state in Eq. (19) is an equal-amplitude superposition of $\eta^{\prime}=+1$ and $\eta^{\prime}=-1$. In this case, the normalized projection onto each value of $\eta^{\prime}$ is
@@
$\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix},\eta^{\prime}\biggr{\rangle}\\ \equiv\sqrt{2}\,\Pi_{\eta^{\prime}}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle},$ (23)
@@
$\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1}(+),\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}\Bigr{)}J_{2}(-)\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle}=(-1)^{2j_{1}}\sqrt{\tilde{J}_{1}\tilde{J}_{2}}\\ \times\sum_{j^{\prime},S_{Q}}\sqrt{\tilde{\jmath}^{\prime}\tilde{S}_{Q}}\begin{Bmatrix}\tfrac{1}{2}&\tfrac{1}{2}&S_{Q}\\ j_{1}&j_{2}&j^{\prime}\\ J_{1}&J_{2}&S\\ \end{Bmatrix}\biggl{|}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+}(+),\tfrac{1}{2}^{-}(-)\Bigr{)}S_{Q},\Bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\Bigr{)}j^{\prime}\Bigr{]}\begin{matrix}S\\[-5.0pt] m_{S}\end{matrix}\biggr{\rangle},$ (24)
where $\tilde{J}=2J+1$ and $\left\{\begin{smallmatrix}j_{1}&j_{2}&j_{3}\\ j_{4}&j_{5}&j_{6}\\ j_{7}&j_{8}&j_{9}\end{smallmatrix}\right\}$ is a Wigner 9-$j$ symbol. The sign $(-1)^{2j_{1}}$ is due to changing the order of the light-QCD operator for the static hadron $j_{1}^{\pi_{1}}$ and the fermionic operator for the heavy antiquark state $\tfrac{1}{2}^{-}$.
III Decays of a Hidden-Heavy Hadron
In this section, we determine the general form of the coupling potential between a confining B-O potential and a hadron-pair potential. We use it to derive selection rules for decays of a hidden-heavy hadron and analytic results for relative partial decay rates.